1 mole of carbon-12 contains 6.022*10^23 carbon-12 atoms with a mass of 12g.
What is the mass of 1 carbon-12 atom using scientific notation?
What is the mass of carbon-12 atoms in 20.0g of carbon-12?
Calculate the number of moles of carbon-12 atoms in 20.0g of carbon-12

Answer by Theo(13342)

i'm not quite sure what you're asking here.

you state that 1 mole of carbon 12 has a mass of 12 grams and contains 6.022 * 10^23 atoms of carbon 12.

this means that 6.022 * 10^23 atoms of carbon 12 has a mass of 12 grams.

you can solve the the mass of one carbon 12 atom by dividing 12 by 6.022 * 10^23.

since 12 is equivalent to 12 * 10^0, you would then divide 12 * 10^0 by 6.022 * 10^23 which will result in 12/6.022 * 10^0/10^23 which will then be equal to 1.992693457 * 10^-23 grams.

that should answer your first question which was:

what is the mass of 1 carbon 12 atom.

the solution is that one carbon 12 atom has a mass of 1.992693457 * 10^-23 grams.

your second question was:

What is the mass of carbon-12 atoms in 20.0g of carbon-12?

the solution to that should be 20 grams.

you might have asked how many atoms of carbon 12 in 20 grams of carbon 12.

that would then be 20/12 * 6.022 * 10^23 = 10.03666667 * 10^23 atoms.

if you multiply that by the mass per gram of carbon 12, you should get 20.

i did and i got 20 grams, as i should have.

your last question asks how many moles are in 20 grams of carbon 12.

if 1 mole weights 12 grams, then take 20 and divide it by 12 and you will get 20/12 = 1.6666666666.... moles, which is equal to 1 and 2/3 moles.

i think this is right.
check it out and get back to me if you don't agree, or if the solutions that you're supposed to get don't agree.